A particle P is let fall vertically from a height to the ground. Another particle Q is projected vertically down 3s later with a velocity of 3m/s^s from the same height. If both particles reached the ground simultaneously. Calculate:

a. time by each particle to reach the ground

b. distance travelled by each particle to reach the ground.
1
Expert's answer
2020-04-08T10:44:14-0400

Let time taken by first particle be "t."

If second particle is projected with velocity"=3\frac{m}{sec}"

For first particle, "d=\frac{1}{2}gt^2"

For second particle, "d=3(t-3)+\frac{1}{2}g(t-3)^2"

Equating both the equations, we get

"\frac{1}{2}gt^2=3(t-3)+\frac{1}{2}g(t-3)^2"

"t=\frac{4}{3}sec"

Putting "t=\frac{4}{3}sec" in any of the equations, we get "d=\frac{80}{9}m"